basha's floating time is 15 minutes longer than zoey's. Together they float a total of 75 minutes. how long did zoey float?
Answer by mananth(16946)

Zoey's floating time = x
Basha's floating time = x+15
x+x+15=75
2x+15=75-15
2x=60
/2
x= 30 minutes Joey's time
45 minutes is Basha's time